a tank has two outlets. Pipe A fills the take in 3.5 hours. Pipes A and B fill the same tank in 3 hours. how long will it take for pipe B to fill the tank?
Answer by richwmiller(17219)

3/3.5+3/b=1
solve for b
b=21
It should have been obvious that pipe b doesn't do much work since pipe a can do the whole job in 3.5 hours and together they only save 30 minutes.
pipe a does 6/7 of the job and pipe b does 3/21=1/7